OpenStudy (anonymous):

Solve using he square root method. 9x^2 - 25 = 0

OpenStudy (cwrw238):

9x^2 - 25 = 0 9x^2 = 25 take sprouts of both sides 3x = 5

OpenStudy (cwrw238):

lol - sprouts - i meant sq roots!!

OpenStudy (anonymous):

9x^2-25=0 =>9x^2=25 =>x^2=25/9. so x=5/3 or -5/3

OpenStudy (anonymous):

thanks